# The SRG/eROSITA All-Sky Survey: X-ray beacons at late cosmic dawn

### v1

@@ Abstract
Context. The Spectrum Roentgen Gamma (SRG)/eROSITA All-Sky Survey (eRASS) is expected to contain $\sim 100$ quasars that emitted their light when the universe was less than a billion years old, i.e. at $z>5.6$. By selection, these quasars populate the bright end of the AGN X-ray luminosity function and their count offers a powerful demographic diagnostic of the parent super-massive black hole population.
@@
Results. We report the discovery of five new quasars in the redshift range $6.15.6<z<$ 6.1. Two of these quasars are detected in eRASS and are by selection X-ray ultra-luminous. These quasars are also detected at radio frequencies. The first one is a broad absorption line quasar which shows significant, order-of-magnitude X-ray dimming over 3.5 years, i.e. about 6 months in the quasar rest frame. The second radio-detected quasar is a jetted source with compact morphology. We show that a blazar configuration is likely for this source, making it the second most distant blazar known to date.
Conclusions. With our pilot study, we demonstrate the power of eROSITA as a discovery machine for luminous quasars in the epoch of reionization. The X-ray emission of the two eROSITA detected quasars are likely to be driven by different high-energetic emission mechanisms a diversity which will be further explored in a future systematic full-hemisphere survey.
@@ 1 Introduction
The study of accreting super-massive black holes (SMBHs) in the first gigayear after the Big Bang reshapes our understanding of black hole formation and growth, with potentially deep implications on the evolution of their host galaxies (e.g. Habouzit et al. 2022). With the advent of deep infrared (IR) space-based imaging, as well as multi-object and slitless spectroscopy with the James Webb Space Telescope (JWST), the redshift frontier of accreting SMBHs has been pushed far into cosmic dawn, to $z>8$ (e.g. Harikane et al. 2023; Larson et al. 2023; Bunker et al. 2023; Maiolino et al. 2023). Combining low-resolution prism spectroscopy with JWST/NIRSpec and with the high-spatial-resolution of the Chandra X-ray space telescope, Goulding et al. (2023) have recently discovered a (lensed) SMBH-powered active galactic nucleus (AGN) at z=10.1, corresponding to a lookback-time of $\sim$13.2 Gyr. The most extreme SMBHs arguably reside in the cores of quasars, luminous persistent panchromatic beacons, that are radiating up to $L_{\mathrm{bol}}\sim 10^{48}\mathrm{erg\,s^{-1}}$ at cosmic times reaching back deep into the epoch of reionization. While the less massive SMBHs discovered by JWST may grant us access to crisp snapshots of earlier evolutionary stages of AGN and their hosts, the mere existence of matured quasars and their central $\sim 10^{9}\,M_{\odot}$ SMBHs in the first gigayear of the universe (up to $z\sim 7.5$, e.g. Bañados et al. 2018; Wang et al. 2021b; Yang et al. 2021) poses a puzzling progenitor problem: How did the seeds of black holes evolve to such massive systems in the cosmic timespan of a few hundred million years? Even under the unlikely assumption of sustained Eddington-limited accretion, well-understood stellar seeding models based on the collapse of early Pop III stars, struggle to explain the formation of the most massive SMBHs found in the cores of $z>6.5$ quasars (for reviews see Inayoshi et al. 2020; Fan et al. 2023). Due to the steep decline of the AGN population with luminosity at high redshift, quasars with $L_{\mathrm{bol}}\sim 10^{48}\,\mathrm{erg\,s^{-1}}$ are rare in the first gigayear (Jiang et al. 2016; Vito et al. 2018; Kulkarni et al. 2019; Matsuoka et al. 2023; Schindler et al. 2023). The low space-density ($\sim 1\;\mathrm{Gpc^{-3}}$) of these objects implies that wide-field surveys, sampling large cosmological volumes, are needed for their search. The optical and near-IR (NIR) Euclid Wide Survey will enable the discovery of numerous $z>7.5$ quasars (Euclid Collaboration et al. 2019) in the coming years.
@@
At slightly lower redshifts, the most luminous quasars at $z\sim 6$, impose less stringent constraints on black hole seed models but set a strict lower limit on the black hole population in the early universe and their accretion history (Vito et al. 2018; Wolf et al. 2021, 2023; Barlow-Hall et al. 2022). These rare AGN populating the extreme end of the luminosity and mass scale therefore constitute powerful demographic probes (e.g. Belladitta et al. 2020). So far about $\sim 475$ reionisation-era quasars have been discovered in wide optical and IR surveys (Fan et al. 2023, and references therein). Selection methods are based on the blue-band dropout photometry imprinted by neutral hydrogen absorption by the inter-Galactic medium (IGM Madau & Rees 2000). Quasar searches in imaging data are challenging as their optical and IR colours are similar to high-surface-density Galactic contaminants: late-type stars (M-, L- and T-dwarfs, e.g. Caballero et al. 2008). M-dwarfs, the most numerous stars in the galaxy (Chabrier 2001), emit in the X-rays due to coronal activity and flares (e.g. Briggs & Pye 2004; Rutledge et al. 2000; Stelzer et al. 2013; Magaudda et al. 2022). However, at $z>5.7$, the bulk of galactic colour-contaminants has spectral type SpT M7 or later (Yang et al. 2017). These objects, dubbed ultra-cool dwarfs (UCDs), are intrinsically X-ray faint in their quiescent state. Stelzer et al. (2022) have investigated the X-ray emission of UCDs in the first pass all-sky survey (eRASS1111The number 1 refers to the number of all-sky surveys Merloni et al. 2024) of the extended Röntgen Survey with an Imaging Telescope Array (eROSITA, Predehl et al. 2021) onboard the Spektrum-Röntgen-Gamma (SRG) mission (Sunyaev et al. 2021). Out of 4754 spectroscopically confirmed UCDs (14914 candidates) in their sample, 21 (37) are associated with an eRASS1 source, i.e. less than $0.3\%$ of the input catalogue. Stelzer et al. (2022) show that the eRASS1-detected UCDs have large X-ray to bolometric luminosity ratios (log $L_{\mathrm{X}}/L_{\mathrm{bol}}\sim-3$), indicative of flaring activity. They conclude that the eRASS1 survey is only sensitive to flaring or the closest quiescent UCDs ($<20\,\mathrm{pc}$). We note that a volume complete X-ray survey of nearby M-dwarfs ($<10\,\mathrm{pc}$) based on eRASS:4 was presented by Caramazza et al. (2023).
@@
The detection of X-rays in eROSITA can therefore significantly reduce the degeneracy between UCDs and high-redshift quasars using the ubiquitous, strong coronal and jet X-ray emission of the latter. In this work, we present a pilot $z\sim 6$ quasar discovery program in eRASS1 through eRASS:4. The sensitivity of eROSITA to the X-ray emission of $z\sim 6$ quasars has already been demonstrated by Wolf et al. (2021): we expect eROSITA to detect $\sim 100$ of the most X-ray luminous quasars ($\mathrm{log}\,L_{\mathrm{X}}>45.5$), at the end of the mission (eRASS:8). In this luminosity-redshift regime, the strong X-ray emission arises from a range of high-energetic physical effects including Doppler boosting of non-thermal jet emission (e.g. Belladitta et al. 2020), inverse Comptonization of cosmic microwave background photons (iC-CMB, e.g. Ighina et al. 2019; Medvedev et al. 2020) and rapid accretion (e.g. Wolf et al. 2023; Zappacosta et al. 2023). eROSITA therefore probes a diversity of X-ray boosting mechanisms whose contributions are yet to be disentangled (e.g. Ighina et al. 2021). Our selection pipeline combines optical, IR and X-ray information. Similar panchromatic searches, spanning multiple orders of magnitudes in frequency, have recently been successfully performed by Gloudemans et al. (2022); Ighina et al. (2023). In these works, the authors based their selection on the detection of jets at radio frequencies with the LOw-Frequency ARray (LOFAR, van Haarlem et al. 2013) and the Australian square kilometre array pathfinder (ASKAP, Hotan et al. 2021) instead of accretion signatures in X-rays. We report the discovery of 5 new $z>5.6$ quasars in the joint Dark Energy Survey data release 2 (DES DR2222In this work, DES always refers to the second data release of the survey., Dark Energy Survey Collaboration et al. 2016) footprint, two of which are robustly detected in the eRASS catalogs. The newly discovered eRASS quasars are rare and extremely energetic phenomenons at high redshift, which we showcase in this work: an X-ray variable broad absorption line (BAL) quasar and a blazar, i.e. a quasar for which the jet axis closely aligns with the line of sight.
@@
This paper is organised as follows. In Section 2 we present the optical, IR, and X-ray based quasar selection pipeline. Discoveries through spectroscopic confirmation are presented in Section 3. In Section 4, we present the properties of the two newly discovered eRASS quasars. We discuss the implication of these discoveries and present our conclusions in Section 5. Throughout this work, magnitudes are given in the AB reference system. Unless specified otherwise, uncertainties are given at the $1\sigma$ level. Distance calculations assume a flat $\Lambda$CDM cosmology (Planck Collaboration et al. 2020).
@@ 2.1 Candidate selection in optical/IR
We have selected $z>5.5$ quasar candidates using optical data from DES. In addition, NIR and mid-IR (MIR) photometry from the VISTA Hemisphere Survey (VHS, McMahon et al. 2013) Data Release 5 and CatWISE2020 (Marocco et al. 2021) complement the DES data in our selection pipeline. The combination of optical and infrared broad-band photometry covers a significant portion of the $z>5.6$ quasar restframe UV/optical SED and its efficiency for quasar discovery has previously been demonstrated (e.g. Reed et al. 2019).
@@ 2.1.1 Colour pre-selection
In addition, we always impose signal-to-noise thresholds on the z-band photometry (S/N ¿ 10) and on the Y-band photometry (S/N ¿ 5). We further require no detection in the g band (S/N ¡ 3). These colour cuts are summarized in Fig. 1. The colour-cuts are more relaxed than the ones used in previous $z\sim 6$ quasar searches based on colour criteria (e.g. Fan et al. 2001; Mortlock et al. 2011; Reed et al. 2015, 2019; Banados et al. 2022; Matsuoka et al. 2022), as the aim here is merely to pre-select dropout objects. We query a sub-sample of 511,654 candidates.
Figure 1: Colour-selection diagram. The hashed and grey regions correspond to the colour-space from which we pre-selected high-redshift quasar candidates for the present pilot survey. Gaussian Kernel density estimates of DES detected M-dwarfs (Best et al. 2018) and L/T-dwarfs (Liu et al. 2002; Hawley et al. 2002; Cruz et al. 2003, 2007; Chiu et al. 2008; Reid et al. 2008; Schmidt et al. 2010; Burningham et al. 2010; Kirkpatrick et al. 2010, 2011; Wahhaj et al. 2011; Burningham et al. 2013; Day-Jones et al. 2013; Marocco et al. 2013, 2014, 2015; Best et al. 2015; Cardoso et al. 2015; Tinney et al. 2018) are displayed as shaded green and shaded blue areas, respectively. DES detected, spectroscopically confirmed $z>5.6$ quasars compiled in Fan et al. (2023) and additional sources from Yang et al. (2024) are shown as salmon red circles. 1D colour distributions of the samples from the literature are shown in the marginal histograms. Magenta stars indicate new quasars discovered in this work. Four of these discoveries are located at the edge of the grey selection box, close to the contamination region of brown dwarfs.
@@ 2.1.2 SED template fitting
As the main candidate filtering step, we opted for spectral energy distribution (SED) template fitting and photometric redshift computation. Again the DES magnitudes in the bands g,r ,i,z and Y are derived using the Kron aperture radius (mag_auto). In addition to the DES photometry, we obtained magnitudes in the NIR and MIR wavebands, by astrometrically cross-matching the main DES catalogue with the VHS and CatWISE2020 source catalogues within $<1^{\prime\prime}$. By construction, we therefore only consider sources that are simultaneously detected in DES, VHS and CatWISE2020, reducing our sample to 220,221 sources. The combination of the relatively deep DES optical photometry with these shallower IR surveys biases our selection towards very red, potentially dusty, sources. Using the photometric code Le PHARE (v2.2, Arnouts & Ilbert 2011), we performed a $\chi^{2}$-based template fitting analysis on the collected photometry of the colour pre-selected sample (see Section 2.1.1). All magnitudes are directly read from the catalogues. For the VHS bands Y, J, H and Ks we have used the default point source aperture corrected magnitudes (aperture 3, 2” in diameter). Finally, for CatWISE2020 MIR magnitude measurements, we have selected profile-fit photometry corrected for proper motion in the bands W1 and W2 (w1mpropm and w2mpropm).
@@
Le PHARE was supplemented with a custom template library for AGN and galaxies already used for the eROSITA Final Equatorial-Depth Survey by Salvato et al. (2022). The original AGN templates used in this work are from Polletta et al. (2007); Salvato et al. (2009); Brown et al. (2019). We also used mixed AGN-galaxy templates from Salvato et al. (2009) and Ananna et al. (2017), based on inactive galaxy templates from Noll et al. (2004). Due to their steep red spectra, UCDs of the spectral classes M, L and T are known to be the most prominent contaminants in photometric searches of neutral-IGM absorbed, high-redshift quasars, as both classes of objects appear as dropouts in optical bands. The spectral slope in the IR bands can help disentangle the typical quasar power-law accretion disk emission from the Rayleigh-Jeans tails of brown dwarf spectra (for a review and quasar searches and contamination issues see Fan et al. 2023). In addition to stellar templates from Pickles (1998) and Bohlin et al. (1995), we account for the spectral diversity of these Galactic colour-contaminants, by including the stellar templates provided by Chabrier et al. (2000); Knapp et al. (2004); Golimowski et al. (2004); Chiu et al. (2006); Bayo et al. (2011).
We tested our SED fitting analysis on all spectroscopically confirmed quasars at $z>5.6$ listed by Fan et al. (2023), with detections within $<1^{\prime\prime}$ in DES, VHS and CatWISE2020. This sample contains 35 quasars (initial discoveries by Mortlock et al. 2009; Bañados et al. 2016; Reed et al. 2017; Yang et al. 2017; Chehade et al. 2018; Decarli et al. 2018; Reed et al. 2019; Shen et al. 2019; Eilers et al. 2020; Venemans et al. 2020; Yang et al. 2021; Matsuoka et al. 2022; Banados et al. 2022). The Le PHARE photo-z calculating sub-routine zphota returns the $\chi^{2}$ values for the best-fit AGN and stellar templates. Out of the 35 currently confirmed DES, VHS and CatWISE2020 z¿5.5 quasars, 34 are assigned a photometric redshift $z_{phot}>5.5$. 32 of these confirmed quasars have been fit AGN models with reduced $\chi_{\rm best}^{2}$ smaller than the reduced $\chi_{\rm star}^{2}$ values for the stellar models, i.e. the AGN model is favoured. The comparison of our photometric estimates of the redshift of these quasars and their actual spectroscopic redshifts is presented in Fig. 2.
@@
Figure 2: Le PHARE test run on previously spectroscopically confirmed $z>5.5$ quasars. The estimated photometric redshift $z_{\rm phot}$ is compared to the true spectroscopic value $z_{\rm spec}$ in the upper panel. The ratio of $\chi^{2}$ of the stellar and the AGN or galaxy template are shown in the lower panel. In only two cases, the stellar solution is preferred.
@@
Figure 3: Optical colour-plane of selected of 2605 $z_{\mathrm{phot}}>5.6$ quasar candidates from our colour-cuts and SED fitting. The candidates (circles) are colour-coded according to their photometric redshift. We show redshift-colour tracks for Le PHARE templates: a typical quasar (pl_QSO_DR2_029_t0, an SDSS-DR5 composite modified by Salvato et al. 2009) and a Seyfert II template (Sey2_template_norm, Polletta et al. 2007), with increasing extinction values (0.0, 0.1, 0.2) displayed by increasing transparency. Squares indicate selected $z_{\mathrm{phot}}>5.6$ candidates that are already confirmed in literature (references in Fan et al. 2023). The five quasars discovered in this work are shown a magenta stars.

### v2

@@ Abstract
Context. The Spectrum Roentgen Gamma (SRG)/eROSITA All-Sky Survey (eRASS) is expected to contain $\sim 100$ quasars that emitted their light when the universe was less than a billion years old, i.e. at $z>5.6$. By selection, these quasars populate the bright end of the AGN X-ray luminosity function, and their space density offers a powerful demographic diagnostic of the parent super-massive black hole population.
@@
Results. We report the discovery of five new quasars in the redshift range $6.15.6<z<$ 6.1. Two of these quasars are detected in eRASS and are, therefore, X-ray ultra-luminous by selection. We also report the detection of these quasars at radio frequencies. The first one is a broad absorption line quasar, which shows significant, order-of-magnitude X-ray dimming over 3.5 years, i.e. about six months in the quasar rest frame. The second X-ray detected quasar is a jetted source with compact morphology. We show that a blazar configuration is likely for this source, making it one of the most distant blazar known to date.
Conclusions. With our pilot study, we demonstrate the power of eROSITA as a discovery machine for luminous quasars in the epoch of reionization. The X-ray emission of the two eROSITA detected quasars are likely to be driven by different high-energetic emission mechanisms, a diversity which we will further explore in a future systematic full-hemisphere survey.
@@ 1 Introduction
The study of accreting super-massive black holes (SMBHs) in the first gigayear after the Big Bang reshapes our understanding of black hole formation and growth, with potentially profound implications on the evolution of their host galaxies (e.g. Habouzit et al. 2022). With the advent of deep infrared (IR) space-based imaging, as well as multi-object and slitless spectroscopy with the James Webb Space Telescope (JWST), the redshift frontier of accreting SMBHs has been pushed far into cosmic dawn, to $z>8$ (e.g. Harikane et al. 2023; Larson et al. 2023; Bunker et al. 2023; Maiolino et al. 2023). Combining low-resolution prism spectroscopy with JWST/NIRSpec and with the high-spatial-resolution of the Chandra X-ray space telescope, Goulding et al. (2023) have recently discovered a (lensed) SMBH-powered active galactic nucleus (AGN) at z=10.1, corresponding to a lookback-time of $\sim$13.2 Gyr. The most extreme SMBHs arguably reside in the cores of quasars, luminous persistent panchromatic beacons, that are radiating up to $L_{\mathrm{bol}}\sim 10^{48}\mathrm{erg\,s^{-1}}$ at cosmic times reaching back deep into the epoch of reionization. While the less massive SMBHs discovered by JWST may grant us access to crisp snapshots of earlier evolutionary stages of AGN and their hosts, the mere existence of matured quasars and their central $\sim 10^{9}\,M_{\odot}$ SMBHs in the first gigayear of the universe (up to $z\sim 7.5$, e.g. Bañados et al. 2018; Wang et al. 2021b; Yang et al. 2021) poses a puzzling progenitor problem: How did the seeds of black holes evolve to such massive systems in the cosmic timespan of a few hundred million years? Even under the unlikely assumption of sustained Eddington-limited accretion, well-understood stellar seeding models based on the collapse of early Pop III stars struggle to explain the formation of the most massive SMBHs found in the cores of $z>6.5$ quasars (for reviews see Inayoshi et al. 2020; Fan et al. 2023). Due to the steep decline of the AGN population with luminosity at high redshift, quasars with $L_{\mathrm{bol}}\sim 10^{48}\,\mathrm{erg\,s^{-1}}$ are rare in the first gigayear (Jiang et al. 2016; Vito et al. 2018; Kulkarni et al. 2019; Matsuoka et al. 2023; Schindler et al. 2023). The low space-density ($\sim 1\;\mathrm{Gpc^{-3}}$) of these objects implies that wide-field surveys, sampling large cosmological volumes, are needed for their search. The optical and near-IR (NIR) Euclid Wide Survey will enable the discovery of numerous $z>7.5$ quasars (Euclid Collaboration et al. 2019) in the coming years.
@@
At slightly lower redshifts, the most luminous quasars at $z\sim 6$, impose less stringent constraints on black hole seed models but set a strict lower limit on the black hole population in the early universe and their accretion history (Vito et al. 2018; Wolf et al. 2021, 2023; Barlow-Hall et al. 2023). These rare AGN populating the extreme end of the luminosity and mass scale constitute powerful demographic probes (e.g. Belladitta et al. 2020). So far, about $\sim 475$ reionisation-era quasars have been discovered in wide optical and IR surveys (Fan et al. 2023, and references therein). Selection methods are based on the blue-band dropout photometry imprinted by neutral hydrogen absorption by the intergalactic medium (IGM Madau & Rees 2000). Quasar searches in imaging data are challenging as their optical and IR colours are similar to high-surface-density Galactic contaminants: late-type stars (M-, L- and T-dwarfs, e.g. Caballero et al. 2008). M-dwarfs, the most numerous stars in the galaxy (Chabrier 2001), emit in the X-rays due to coronal activity and flares (e.g. Briggs & Pye 2004; Rutledge et al. 2000; Stelzer et al. 2013; Magaudda et al. 2022). However, at $z>5.7$, the bulk of galactic colour-contaminants has spectral type SpT M7 or later (Yang et al. 2017). These objects, dubbed ultra-cool dwarfs (UCDs), are intrinsically X-ray faint in their quiescent state. Stelzer et al. (2022) have investigated the X-ray emission of UCDs in the first pass all-sky survey (eRASS1111The number 1 refers to the number of all-sky surveys Merloni et al. 2024) of the extended Röntgen Survey with an Imaging Telescope Array (eROSITA, Predehl et al. 2021) onboard the Spektrum-Röntgen-Gamma (SRG) mission (Sunyaev et al. 2021). Out of 4754 spectroscopically confirmed UCDs (14914 candidates) in their sample, 21 (37) are associated with an eRASS1 source, i.e. less than $0.3\%$ of the input catalogue. Stelzer et al. (2022) show that the eRASS1-detected UCDs have large X-ray to bolometric luminosity ratios (log $L_{\mathrm{X}}/L_{\mathrm{bol}}\sim-3$), indicative of flaring activity. They conclude that the eRASS1 survey is only sensitive to flaring or the closest quiescent UCDs ($<20\,\mathrm{pc}$). We note that a volume complete X-ray survey of nearby M-dwarfs ($<10\,\mathrm{pc}$) based on eRASS:4 was presented by Caramazza et al. (2023).
@@
Thus, the detection of X-rays in eROSITA can significantly reduce the degeneracy between UCDs and high-redshift quasars using the latter’s ubiquitous, strong coronal and jet X-ray emission. In this work, we present a pilot $z\sim 6$ quasar discovery program in eRASS1 through eRASS:4. Wolf et al. (2021) have already demonstrated the sensitivity of eROSITA to the X-ray emission of $z\sim 6$ quasars: we expect eROSITA to detect $\sim 100$ of the most X-ray luminous quasars ($\mathrm{log}\,L_{\mathrm{X}}>45.5$), at the end of the mission (eRASS:8). In this luminosity-redshift regime, the strong X-ray emission arises from a range of high-energetic physical effects including Doppler boosting of non-thermal jet emission (e.g. Belladitta et al. 2020), inverse Comptonization of cosmic microwave background photons (iC-CMB, e.g. Ighina et al. 2019; Medvedev et al. 2020) and rapid accretion (e.g. Wolf et al. 2023; Zappacosta et al. 2023). eROSITA thus probes a diversity of X-ray boosting mechanisms whose contributions are yet to be disentangled (e.g. Ighina et al. 2021). Our selection pipeline combines optical, IR and X-ray information. Similar panchromatic searches, spanning multiple orders of magnitudes in frequency, have recently been successfully performed by Gloudemans et al. (2022); Ighina et al. (2023). In these works, the authors based their selection on the detection of jets at radio frequencies with the LOw-Frequency ARray (LOFAR, van Haarlem et al. 2013) and the Australian square kilometre array pathfinder (ASKAP, Hotan et al. 2021) instead of accretion signatures in X-rays. We report the discovery of 5 new $z>5.6$ quasars in the joint Dark Energy Survey data release 2 (DES DR2222In this work, DES always refers to the second data release of the survey., Dark Energy Survey Collaboration et al. 2016) footprint, two of which are robustly detected in the eRASS catalogues. The newly discovered eRASS quasars are rare and highly energetic phenomenons at high redshift, which we showcase in this work: an X-ray variable broad absorption line (BAL) quasar and a blazar, i.e. a quasar for which the jet axis closely aligns with the line of sight.
@@
This paper is organised as follows. In Section 2 we present the optical, IR, and X-ray based quasar selection pipeline. Discoveries through spectroscopic confirmation are presented in Section 3. In Section 4, we present the properties of the two newly discovered eRASS quasars. We discuss the implication of these discoveries and present our conclusions in Section 5. Throughout this work, magnitudes are given in the AB reference system. Unless specified otherwise, uncertainties are given at the $1\sigma$ level. Distance calculations assume a flat $\Lambda$CDM cosmology (Planck Collaboration et al. 2020, $H_{0}=67.66\pm 0.42\,\mathrm{km/(Mpc\,s)}$, $\Omega_{m}=0.3111\pm 0.0056$, $\Omega_{\Lambda}=0.6889\pm 0.0056$, $\Omega_{K}=0.001\pm 0.002$)
@@ 2.1 Candidate selection in optical/IR
We have selected $z>5.6$ quasar candidates using optical data from DES. In addition, NIR and mid-IR (MIR) photometry from the VISTA Hemisphere Survey (VHS, McMahon et al. 2013) Data Release 5 and CatWISE2020 (Marocco et al. 2021) complement the DES data in our selection pipeline. The combination of optical and infrared broad-band photometry covers a significant portion of the $z>5.6$ quasar restframe UV/optical SED, and its efficiency for quasar discovery has previously been demonstrated (e.g. Reed et al. 2019).
@@ 2.1.1 Colour pre-selection
In addition, we always impose signal-to-noise thresholds on the z-band photometry (S/N ¿ 10) and on the Y-band photometry (S/N ¿ 5). We further require no detection in the g band (S/N ¡ 3). These colour cuts are summarized in Fig. 1. The colour cuts are more relaxed than the ones used in previous $z\sim 6$ quasar searches based on colour criteria (e.g. Fan et al. 2001; Mortlock et al. 2011; Reed et al. 2015, 2019; Bañados et al. 2023; Matsuoka et al. 2022), as the aim here is merely to pre-select dropout objects. We query a sub-sample of 511,654 candidates.
Figure 1: Colour-selection diagram. The hashed and grey regions correspond to the colour space from which we pre-selected high-redshift quasar candidates for the present pilot survey. Gaussian Kernel density estimates of DES detected M-dwarfs (Best et al. 2018) and L/T-dwarfs (Liu et al. 2002; Hawley et al. 2002; Cruz et al. 2003, 2007; Chiu et al. 2008; Reid et al. 2008; Schmidt et al. 2010; Burningham et al. 2010; Kirkpatrick et al. 2010, 2011; Wahhaj et al. 2011; Burningham et al. 2013; Day-Jones et al. 2013; Marocco et al. 2013, 2014, 2015; Best et al. 2015; Cardoso et al. 2015; Tinney et al. 2018) are displayed as shaded green and shaded blue areas, respectively. DES detected, spectroscopically confirmed $z>5.6$ quasars compiled in Fan et al. (2023) and additional sources from Yang et al. (2024) are shown as salmon red circles. 1D colour distributions of the samples from the literature are shown in the marginal histograms. Magenta stars indicate new quasars discovered in this work. Four of these discoveries are located at the edge of the grey selection box, close to the contamination region of brown dwarfs.
@@ 2.1.2 SED template fitting
As the main candidate filtering step, we opted for spectral energy distribution (SED) template fitting and photometric redshift computation. Again the DES magnitudes in the bands g,r ,i,z and Y are derived using the Kron aperture radius (mag_auto). In addition to the DES photometry, we obtained magnitudes in the NIR and MIR wavebands by astrometrically cross-matching the main DES catalogue with the VHS and CatWISE2020 source catalogues within $<1^{\prime\prime}$. By construction, we only consider sources simultaneously detected in DES, VHS and CatWISE2020, reducing our sample to 220,221 sources. Combining the relatively deep DES optical photometry with these shallower IR surveys biases our selection towards very red, potentially dusty sources. Using the photometric code Le PHARE (v2.2, Arnouts & Ilbert 2011), we performed a $\chi^{2}$-based template fitting analysis on the collected photometry of the colour pre-selected sample (see Section 2.1.1). All magnitudes are directly read from the catalogues. For the VHS bands Y, J, H and Ks we have used the default point source aperture corrected magnitudes (aperture 3, 2” in diameter). Finally, for CatWISE2020 MIR magnitude measurements, we have selected profile-fit photometry corrected for proper motion in the bands W1 and W2 (w1mpropm and w2mpropm).
@@
Le PHARE was supplemented with a custom template library for AGN and galaxies already used for the eROSITA Final Equatorial-Depth Survey by Salvato et al. (2022). The original AGN templates used in this work are from Polletta et al. (2007); Salvato et al. (2009); Brown et al. (2019). We also used mixed AGN-galaxy templates from Salvato et al. (2009) and Ananna et al. (2017), based on inactive galaxy templates from Noll et al. (2004). Due to their steep red spectra, UCDs of the spectral classes M, L and T are known to be the most prominent contaminants in photometric searches of neutral-IGM absorbed, high-redshift quasars, as both classes of objects appear as dropouts in optical bands. The spectral slope in the IR bands can help disentangle the typical quasar power-law accretion disk emission from the Rayleigh-Jeans tails of brown dwarf spectra (for a review and quasar searches and contamination issues see Fan et al. 2023). In addition to stellar templates from Pickles (1998) and Bohlin et al. (1995), we account for the spectral diversity of these Galactic colour-contaminants by including the stellar templates provided by Chabrier et al. (2000); Knapp et al. (2004); Golimowski et al. (2004); Chiu et al. (2006); Bayo et al. (2011).
[…]
Figure 2: Le PHARE test run on previously spectroscopically confirmed $z>5.6$ quasars. Upper panel:
@@
The estimated photometric redshift $z_{\rm phot}$ is compared to the true spectroscopic value $z_{\rm spec}$. The grey shaded area indicates the region within which $\Delta z<0.15\times(1+z_{\mathrm{spec}})$ where $\Delta z=\mid z_{\mathrm{spec}}-z_{\mathrm{phot}}\mid$. Lower panel: The ratio of $\chi^{2}$ of the stellar and the AGN or galaxy template are shown in the lower panel. In only two cases, the stellar solution is preferred.
@@
Figure 3: Optical colour-plane of selected of 2605 $z_{\mathrm{phot}}>5.6$ quasar candidates from our colour-cuts and SED fitting. The candidates (circles) are colour-coded according to their photometric redshift. We show redshift-colour tracks for Le PHARE templates: a typical quasar (pl_QSO_DR2_029_t0, an SDSS-DR5 composite modified by Salvato et al. 2009) and a Seyfert II template (Sey2_template_norm, Polletta et al. 2007), with increasing extinction values (0.0, 0.1, 0.2) displayed by increasing transparency. Squares indicate selected $z_{\mathrm{phot}}>5.6$ candidates that are already confirmed in literature (references in Fan et al. 2023). The five quasars discovered in this work are shown as magenta stars.